# Nonsingular black holes and spherically symmetric objects in nonlinear electrodynamics with a scalar field

### v1

@@ Abstract
In general relativity with vector and scalar fields given by the Lagrangian ${\cal L}(F,\phi,X)$, where $F$ is a Maxwell term and $X$ is a kinetic term of the scalar field, we study the linear stability of static and spherically symmetric objects without curvature singularities at their centers. We show that the background solutions are generally described by either purely electrically or magnetically charged objects with a nontrivial scalar-field profile. In theories with the Lagrangian ~$\tilde{{\cal L}}(F)+K(\phi,X)$, which correspond to nonlinear electrodynamics with a k-essence scalar field, angular Laplacian instabilities induced by vector-field perturbations exclude all the regular spherically symmetric solutions including nonsingular black holes. In theories described by the Lagrangian ${\cal L}=X+\mu(\phi)F^{n}$, where $\mu$ is a function of $\phi$ and $n$ is a constant, the absence of angular Laplacian instabilities of spherically symmetric objects requires that $n>1/2$, under which nonsingular black holes with event horizons are not present. However, for some particular ranges of $n$, there are horizonless compact objects with neither ghosts nor Laplacian instabilities in the small-scale limit. In theories given by ${\cal L}=X\kappa(F)$, where $\kappa$ is a function of $F$, regular spherically symmetric objects are prone to Laplacian instabilities either around the center or at spatial infinity. Thus, in our theoretical framework, we do not find any example of linearly stable nonsingular black holes.
@@ I Introduction
Theories we will study in this paper belong to a subclass of scalar-vector-tensor theories with second-order field equations of motion. Since they respect $U(1)$ gauge symmetry, there are one scalar, two transverse vectors, and two tensor polarizations as the propagating DOFs. To derive the stability conditions of those five DOFs, we consider linear perturbations in both odd- and even-parity sectors on the SSS background. We expand the corresponding action up to the second order in perturbations by taking into account both electric and magnetic charges. For the electric case, a similar analysis was performed in Refs. [43, 44] as a subclass of Maxwell-Horndeski theories. Since the linear stability of magnetic SSS objects has not been addressed yet, we will do so in this paper.
@@
After deriving conditions for the absence of ghosts and Laplacian instabilities of the five dynamical DOFs, we will apply them to three subclasses of ${\cal L}(F,\phi,X)$ theories: (i) ~$\tilde{\cal L}(F)+K(\phi,X)$, (ii) ${\cal L}=X+\mu(\phi)F^{n}$, and (iii) ${\cal L}=X\kappa(F)$. We will show that regular SSS objects realized by theories (i) and (iii), which include nonsingular BHs, are excluded by Laplacian instabilities of vector-field perturbations around the origin. In theories (ii), the absence of angular Laplacian instabilities requires the condition $n>1/2$, under which there are no regular BHs with even horizons. Thus, even by extending NED to more general theories with the Lagrangian ${\cal L}(F,\phi,X)$, we do not find even a single example of nonsingular BHs without instabilities. This shows the general difficulty of constructing BHs without singularities in classical field theories. In theories (ii), however, we will show the existence of linearly stable SSS compact objects without event horizons. These regular solutions are present for both electric and magnetic configurations. We will clarify the regions of $n$ in which the regular horizonless compact objects are subject to neither ghosts nor Laplacian instabilities.
@@
This paper is organized as follows. In Sec. II, we derive the SSS background solutions and discuss the properties of them for the electrically and magnetically charged cases. In Sec. III, we expand the action up to quadratic order in perturbations and obtain conditions under which neither ghosts nor Laplacian instabilities are present for five dynamical DOFs. In Sec. IV, we apply the linear stability conditions to theories (i) mentioned above and show that nonsingular SSS objects are prone to angular Laplacian instability. In Sec. V, we show that the absence of angular Laplacian instabilities demands the condition $n>1/2$, under which nonsingular BHs with event horizons do not exist. We also clarify the parameter space of $n$ in which horizonless regular compact objects suffer from neither ghosts nor Laplacian instabilities. In Sec. VI, we show that nonsingular SSS objects in theories (iii) are excluded by Laplacian instabilities either at the origin or at spatial infinity. Sec. VII is devoted to conclusions.
@@ II Field equations on the SSS background
We will consider SSS objects that are regular at $r=0$, including nonsingular BHs with event horizons [2, 3, 4, 5, 6, 7, 8, 9]. In such cases, the Ricci scalar $R$, the squared Ricci tensor $R_{\mu\nu}R^{\mu\nu}$, and the squared Riemann tensor $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ do not diverge at $r=0$. This requires that $h$ and $N$ are expanded around $r=0$, as [45]
@@ V.1 Electric case
Under the linear stability condition (104), $h(r)$ is always positive at any distance $r\geq 0$. Therefore, we do not have the nonsingular BH configuration with an event horizon.
@@
The ADM mass of SSS objects is defined by
@@
In Fig. 2, we plot the ADM mass (129) for seven different values of $n$, with $N_{0}=0.2$. The analytic estimation (126) of $h(r)$ shows that $M$ asymptotically approaches constants for 1/$21/2<n<3/$ 2, whose property can be confirmed in Fig. 2. For $n=3/2$, the ADM mass exhibits the logarithmic growth $M\propto\ln r$ due to the property of Eq. (127). For $n>3/2$, Eq. (128) gives the analytic dependence $M\propto r^{\frac{2n-3}{2n-1}}$. In Fig. 2, we can confirm this power-law growth of $M$ for $n>3/2$ at large distances. Thus, so long as 1/$21/2<n<3/$ 2, the electric SSS configuration without the event horizon can be interpreted as a compact object.
@@ V.2 Magnetic case
Since $h(r)>0$ at any $r\geq 0$ under the Laplacian stability condition $n>1/2$, there are no regular magnetic BHs with event horizons.
@@
In Fig. 6, we show $h$ versus $r/r_{0}$ for $n=1/2,4/5,1,2$ and the limit $n\gg 1$, with the choice $N_{0}=0.1$ in Eq. (120). The theoretical curves are bounded from below by the line $h_{n\to 1/2}(r)$. We have $h(r)>0$ for all the powers $n$ in the range $n>1/2$, showing the absence of nonsingular magnetic BHs with event horizons. The compactness of magnetic SSS objects requires that $n>3/4$. In such cases, we numerically checked the asymptotic behaviors $h-1\propto r^{4}$ for $r\ll r_{0}$ and $h-1\propto r^{-1}$ for $r\gg r_{0}$. We note that the radial dependence of $\phi^{\prime}(r)$ is similar to that of the electric case shown in Fig. 3.
@@ VI.2 Magnetic case
which is positive for $r\geq 0$. Then, the above solution does not correspond to the nonsingular BH with an event horizon.
@@ VII Conclusions
In Sec. V, we analyzed regular solutions for the Lagrangian of the form ${\cal L}=X+\mu(\phi)F^{n}$. For both electric and magnetic SSS objects, we showed that the absence of angular Laplacian instabilities imposes the condition $n>1/2$. Under this inequality, the metric components are positive at any distance $r$, so that there are no regular BHs with event horizons.
@@
Finally, we investigated Lagrangians of the form ${\cal L}=X\kappa(F)$ in Sec. VI. Similar to the previous case, these theories do not support the SSS configurations with event horizons, precluding the existence of nonsingular BHs. Unlike theories discussed in Sec. V, all regular SSS solutions obtained in this framework (including those without horizons) suffer from angular Laplacian instabilities. Specifically, for electric solutions, such instabilities arise around the regular center. In contrast, for magnetic solutions, they manifest at large distances away from the origin.
Thus, we have not found any linearly stable nonsingular BHs in subsets of theories given by the Lagrangian ${\cal L}(F,\phi,X)$. This shows that the construction of regular BHs in the framework of classical field theories is highly complex and challenging. In other words, Penrose’s singularity theorem holds even in scenarios where some of its original assumptions are violated. If this property generally persists for local classical four-dimensional actions, we may need to resort to non-local theories of gravity or higher-dimensional theories (see, e.g., [65, 66, 67, 68, 69, 70]). On the other hand, our analysis in Sec. V has uncovered linearly stable regular SSS solutions without horizons that, in principle, could correspond to physically realizable configurations in nature. The exploration of physical and geometric properties of such stable compact objects without horizons, including the comparisons with boson and Proca stars [71, 72], remains an intriguing avenue for future research.

### v2

@@ Abstract
In general relativity with vector and scalar fields given by the Lagrangian ${\cal L}(F,\phi,X)$, where $F$ is a Maxwell term and $X$ is a kinetic term of the scalar field $\phi$, we study the linear stability of static and spherically symmetric objects without curvature singularities at their centers. We show that the background solutions are generally described by either purely electrically or magnetically charged objects with a nontrivial scalar-field profile. In theories with the Lagrangian ~$\tilde{{\cal L}}(F)+K(\phi,X)$, which correspond to nonlinear electrodynamics with a k-essence scalar field, angular Laplacian instabilities induced by vector-field perturbations exclude all the regular spherically symmetric solutions including nonsingular black holes. In theories described by the Lagrangian ${\cal L}=X+\mu(\phi)F^{n}$, where $\mu$ is a function of $\phi$ and $n$ is a constant, the absence of angular Laplacian instabilities of spherically symmetric objects requires that $n>1/2$, under which nonsingular black holes with apparent horizons are not present. However, for some particular ranges of $n$, there are horizonless compact objects with neither ghosts nor Laplacian instabilities in the small-scale limit. In theories given by ${\cal L}=X\kappa(F)$, where $\kappa$ is a function of $F$, regular spherically symmetric objects are prone to Laplacian instabilities either around the center or at spatial infinity. Thus, in our theoretical framework, we do not find any example of linearly stable nonsingular black holes.
@@ I Introduction
The theories we will study in this paper belong to a subclass of scalar-vector-tensor theories with second-order field equations of motion. Since they respect $U(1)$ gauge symmetry, there are one scalar, two transverse vectors, and two tensor polarizations as the propagating DOFs. To derive the stability conditions of those five DOFs, we consider linear perturbations in both odd- and even-parity sectors on the SSS background. We expand the corresponding action up to the second order in perturbations by taking into account both electric and magnetic charges. For the electric case, a similar analysis was performed in Refs. [43, 44] as a subclass of Maxwell-Horndeski theories. Since the linear stability of magnetic SSS objects has not been addressed yet, we will do so in this paper.
@@
After deriving conditions for the absence of ghosts and Laplacian instabilities of the five dynamical DOFs, we will apply them to three subclasses of ${\cal L}(F,\phi,X)$ theories: (i) ~$\tilde{\cal L}(F)+K(\phi,X)$, (ii) ${\cal L}=X+\mu(\phi)F^{n}$, and (iii) ${\cal L}=X\kappa(F)$. We will show that regular SSS objects realized by theories (i) and (iii), which include nonsingular BHs, are excluded by Laplacian instabilities of vector-field perturbations around the origin. In theories (ii), the absence of angular Laplacian instabilities requires the condition $n>1/2$, under which there are no regular BHs with apparent horizons. Thus, even by extending NED to more general theories with the Lagrangian ${\cal L}(F,\phi,X)$, we do not find even a single example of nonsingular BHs without instabilities.
This shows the general difficulty of constructing BHs without singularities in classical field theories. In theories (ii), however, we will show the existence of linearly stable SSS compact objects without apparent horizons. These regular solutions are present for both electric and magnetic configurations. We will clarify the regions of $n$ in which the regular horizonless compact objects are subject to neither ghosts nor Laplacian instabilities.
@@
This paper is organized as follows. In Sec. II, we derive the SSS background solutions and discuss the properties of them for the electrically and magnetically charged cases. In Sec. III, we expand the action up to quadratic order in perturbations and obtain conditions under which neither ghosts nor Laplacian instabilities are present for five dynamical DOFs. In Sec. IV, we apply the linear stability conditions to theories (i) mentioned above and show that nonsingular SSS objects are prone to angular Laplacian instability. In Sec. V, we show that the absence of angular Laplacian instabilities demands the condition $n>1/2$, under which nonsingular BHs with apparent horizons do not exist. We also clarify the parameter space of $n$ in which horizonless regular compact objects suffer from neither ghosts nor Laplacian instabilities. In Sec. VI, we show that nonsingular SSS objects in theories (iii) are excluded by Laplacian instabilities either at the origin or at spatial infinity. Sec. VII is devoted to conclusions.
@@ II Field equations on the SSS background
We will consider SSS objects that are regular at $r=0$, including nonsingular BHs with apparent horizons [2, 3, 4, 5, 6, 7, 8, 9]. In such cases, the Ricci scalar $R$, the squared Ricci tensor $R_{\mu\nu}R^{\mu\nu}$, and the squared Riemann tensor $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ do not diverge at $r=0$. This requires that $h$ and $N$ are expanded around $r=0$, as [45]
@@ V.1 Electric case
Under the linear stability condition (104), $h(r)$ is always positive at any distance $r\geq 0$. Therefore, we do not have the nonsingular BH configuration with an apparent horizon.
@@
The Arnowitt-Deser-Misner (ADM) mass of SSS objects is defined by
@@
In Fig. 2, we plot the ADM mass (129) for seven different values of $n$, with $N_{0}=0.2$. The analytic estimation (126) of $h(r)$ shows that $M$ asymptotically approaches constants for 1/$21/2<n<3/$ 2, whose property can be confirmed in Fig. 2. For $n=3/2$, the ADM mass exhibits the logarithmic growth $M\propto\ln r$ due to the property of Eq. (127). For $n>3/2$, Eq. (128) gives the analytic dependence $M\propto r^{\frac{2n-3}{2n-1}}$. In Fig. 2, we can confirm this power-law growth of $M$ for $n>3/2$ at large distances. Thus, so long as 1/$21/2<n<3/$ 2, the electric SSS configuration without the apparent horizon can be interpreted as a compact object.
@@ V.2 Magnetic case
Since $h(r)>0$ at any $r\geq 0$ under the Laplacian stability condition $n>1/2$, there are no regular magnetic BHs with apparent horizons.
@@
In Fig. 6, we show $h$ versus $r/r_{0}$ for $n=1/2,4/5,1,2$ and the limit $n\gg 1$, with the choice $N_{0}=0.1$ in Eq. (120). The theoretical curves are bounded from below by the line $h_{n\to 1/2}(r)$. We have $h(r)>0$ for all the powers $n$ in the range $n>1/2$, showing the absence of nonsingular magnetic BHs with apparent horizons. The compactness of magnetic SSS objects requires that $n>3/4$. In such cases, we numerically checked the asymptotic behaviors $h-1\propto r^{4}$ for $r\ll r_{0}$ and $h-1\propto r^{-1}$ for $r\gg r_{0}$. We note that the radial dependence of $\phi^{\prime}(r)$ is similar to that of the electric case shown in Fig. 3.
@@ VI.2 Magnetic case
which is positive for $r\geq 0$. Then, the above solution does not correspond to the nonsingular BH with an apparent horizon.
@@ VII Conclusions
In Sec. V, we analyzed regular solutions for the Lagrangian of the form ${\cal L}=X+\mu(\phi)F^{n}$. For both electric and magnetic SSS objects, we showed that the absence of angular Laplacian instabilities imposes the condition $n>1/2$. Under this inequality, the metric components are positive at any distance $r$, so that there are no regular BHs with apparent horizons.
@@
Finally, we investigated Lagrangians of the form ${\cal L}=X\kappa(F)$ in Sec. VI. Similar to the previous case, these theories do not support the SSS configurations with apparent horizons, precluding the existence of nonsingular BHs. Unlike theories discussed in Sec. V, all regular SSS solutions obtained in this framework (including those without horizons) suffer from angular Laplacian instabilities. Specifically, for electric solutions, such instabilities arise around the regular center. In contrast, for magnetic solutions, they manifest at large distances away from the origin.
To summarize, we have not found any linearly stable nonsingular BHs for theories studied in this paper. To reach this conclusion, we have made three assumptions: (i) the choice of three subsets of ${\cal L}(F,\phi,X)$ theories, (ii) the spherically symmetric and static regular background with asymptotic flatness, and (iii) the existence of the BH apparent horizon. To circumvent the no-go result, one could either modify the theories—by extending the analysis to other subclasses of ${\cal L}(F,\phi,X)$ theories and more general theories—or alter the spacetime background. For instance, rotating regular BH solutions, if they exist, might offer a stable alternative.
Our results show that the construction of regular BHs in the framework of classical field theories is highly complex and challenging. In other words, Penrose’s singularity theorem holds even in scenarios where some of its original assumptions are violated. If this property generally persists for local classical four-dimensional actions, we may need to resort to non-local theories of gravity or higher-dimensional theories (see, e.g., [65, 66, 67, 68, 69, 70]). On the other hand, our analysis in Sec. V has uncovered linearly stable regular SSS solutions without horizons that, in principle, could correspond to physically realizable configurations in nature. The exploration of physical and geometric properties of such stable compact objects without horizons, including the comparisons with boson and Proca stars [71, 72], remains an intriguing avenue for future research.